The values of m for which the expression 2x^2+mxy+3y^2-5y-2 can be expressed as the product of 2 linear factors are+1,-1.


(True or False)

Grade:12

1 Answers

Ramesh V
14 years ago

 

It can be expressed as (2x+y+a)(x+3y+b)=0

which is

2x2 +7xy+3y2+x(2b+a)+y(3a+b)+ab=0

on comparing , we'll have 2b+a=0

              ab= -2

     3a+b=-5

on solving gives

a = -2 and b = 1 and m = 7

so its, (2x+y-2)(x+3y+1)=0

its FALSE